A rope inclined at angle 37degree to the horizontal is used to drag a 50kg block along a level floor with an acceleration of 1m/s². The coefficient of friction between the block and floor is 0.2. What is the tension in the rope?


1
Expert's answer
2021-05-24T15:44:04-0400

Gives

"\theta=37°"

"a=1m\/sec"

m=50kg

"\mu=.2"

We know that

Horizontal components

"f=Tcos37\u00b0"

Vertical components

Tsin37°+N=mg

N=mg-Tsin37°

We can rewrite horizontal components

f=Tcos37°

"\mu N=Tcos37°"

Put N value

"\mu(mg-Tsin37°)=Tcos30°"

"\mu mg=T(\mu sin30°+cos30°)"

"T=\frac{\mu mg}{\mu sin37°+cos37°}"

Put value

"T=\frac{0.2\times50\times9.8}{0.2sin37°+cos37°}"

"T=\frac{98}{0.9189}=106.64" N